OpenStudy (anonymous):

I solved a problem for H and I don't think I'm done... R=P+A(sqrtH-C). I got H=[(R-P)^2/A^2]+C

OpenStudy (lgbasallote):

\[\huge R = P + A \sqrt{H+C}\] \[\huge \implies R - P = A \sqrt{H+C}\] \[\huge \implies \frac{R-P}{A} = \sqrt{H+C}\] \[\huge \implies \left(\frac{R-P}{A} \right)^2 = (\sqrt{H+C})^2\] \[\huge \implies \left(\frac{R-P}{A} \right)^2= H+C\] \[\huge \implies \left(\frac{R-P}{A} \right)^2 - C = H\] you're right...you just needed to "factor out" that exponent

OpenStudy (anonymous):

Oh okay. Thank you!

OpenStudy (lgbasallote):

welcome